OpenStudy (anonymous):

"A hole is punched in a full milk carton, 10 cm below the top. What is the initial velocity of outflow?" This is the only information given.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Bernoulli%27s_principle this should help.

OpenStudy (anonymous):

Bernouli's equation: \[P_1 + \frac{1}{2}\rho v_1^2 + h_1\rho g = P_2 + \frac{1}{2}\rho v_2^2 + h_2\rho g\] With P1 = P2 because they are both exposed to the outside atmosphere v1 approximately = 0 and h2 = 0, that is easy to solve for v2